Falconmaster:

Simplify the given equation. 17x - 6 + 3x - 5 = x + 11 + 4x 20x - 11 = 5x + 11 20x + 1 = 5x + 11 9x = 16x

Mehek:

Do you know how to combine like terms?

Falconmaster:

what in the what ??

Mehek:

Like terms Such as 2y + 3y = 5y they are like terms because both have y So in your problem, you have 17x + 3x Any idea what that would equal?

Falconmaster:

20?

Mehek:

Correct so 20x - 6 - 5 -6 - 5 = ?

Falconmaster:

1

Falconmaster:

so b?

Mehek:

it's not 1

Falconmaster:

oh well then

Falconmaster:

-1?

Falconmaster:

as you can see i am not smart :D

Mehek:

-6 - 5 = 11 So one side is simplified to 20x - 11 and then on the other side you have x + 11 + 4x The like terms are x + 4x which = 5x so 20x - 11 = 11 + 5x

Mehek:

It's ok this isn't a hard concept; you'll catch on quick

AshtynThatDrummerGirl:

Do you still need help or does @Mehek got it?

Falconmaster:

she gots it

Falconmaster:

i might need a 3rd opinion tho

AshtynThatDrummerGirl:

Okay, well I'll stay here and observe and put my input in where I see it needed :)

Falconmaster:

okay

Falconmaster:

An equation _____ has one solution. always sometimes never

Falconmaster:

sometimes?

Mehek:

Correct

Falconmaster:

new post so i can give out medals